solve by substitution or elimination method: 3x-2y=8
-12x+8y=32

Answer by midwood_trail(310)

solve by substitution or elimination method: 3x-2y=8
-12x+8y=32
The first thing to do is to isolate one of the letters.
It is YOUR CHOICE which letter to isolate first.
To isolate means to place the letter of choice ALL ALONE on one side of the equation.
I will isolate y in the equation 3x - 2y = 8.
Subtracting 3x from both sides we get this:
-2y = -3x + 8
The letter y is not alone. See it? It has a friend named coefficient -2.
To get y alone, we now divide both sides by -2.
y = (-3x + 8)/-2
y = (3x/2) - 4
Since we now found y to be (3x/2) - 4, we plug that into -12x + 8y = 32 to find the value of x.
LOOK:
-12x + 8(3x/2) - 4) = 32
-12x + 12x - 32 = 32....STOP!!!!!
If you look carefully, on the left side of the equation you will see that -12x and 12x BOTH CANCEL OUT.
What does this mean?
Since x is gone from the system of equations, then we can STOP RIGHT HERE and say that this particular system of equations in two variables has NO SOLUTION.
This is your final answer: NO SOLUTION.
Did you follow?